Find (i7/24 – 17/36 Find (i7/24 – 17/36 Find (i7/24 – 17/36

Find    (i7/24 – 17/36 
Find    (i7/24 – 17/36 
Find    (i7/24 – 17/36 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Take the LCM of the denominators of the given rational numbers.
LCM of 24 and 36 is 72
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(7/24)= [(7×3)/ (24×3)] = (21/72)
(17/36)= [(17×2)/ (36×2)] = (34/72)
Then,
= (21/72) - (34/72) ...[∵denominator is same in both the rational numbers]
= (21 - 34)/72
= (-13/72)

Thanks